OpenStudy (anonymous):

Which equation could have been used to create this function table? A. X Y 4 6 y = 4x 5 7 7 9 10 12 11 13 B. y = 5x C. y = x + 7 D. y = x + 2

OpenStudy (anonymous):

A

OpenStudy (anonymous):

@snowsurf

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

no guessing. explain why you chose A.

OpenStudy (anonymous):

thought I had it figured out but now I don't know

OpenStudy (anonymous):

I already know the answer. XD

OpenStudy (anonymous):

@countsrnc It is simple math. You should first understand how these equations work to get those coordinates.

OpenStudy (anonymous):

show me

OpenStudy (anonymous):

?

OpenStudy (anonymous):

people tell me different ways and I can't seem to grasp it, I thought I did

OpenStudy (anonymous):

Hint: To find x, you set the y=0. You will get a coordinate (x,0) To find y, you set the x=0. You will get a coordinate (0, y) Apply this method in the problem. By testing it in each equation. But this time, you don't use zeroes.

OpenStudy (anonymous):

@countsrnc You now use the values in the table and test which equation shows the same values in x and y.

OpenStudy (anonymous):

y=6 x=4 difference is 2 which makes it y=x+2

OpenStudy (anonymous):

D

OpenStudy (whpalmer4):

I would look at this by first checking if there is a simple constant difference between x and y. if so, the relationship is of the form \[y = x + c\]where \(c\) is a constant. if that test fails, I then look to see if \{y/x\) always gives the same result, or in other words, \(y/x=c\) in which case the equation is of the form \[y=cx\] if that fails, then I would use the equation for the slope of a line:\[m=\frac{y_2-y_1}{x_2-x_1}\]to find the slope of the line from two points, and then use the point-slope formula to get the equation for the line:\[y-y_1=m(x-x_1)\] between those three tactics, you should be able to determine the equation from a table of linear data like this.